Left Direct Gregorian Technical Observing with IRTC: UT 2009 December 27

Observer: JHill (LBTO), DMiller (Tucson)
Telescope Operator: DHuerta
Instrument Support:
Telescope Support: BCurtis
SW support: (Tucson)
Phone Support: TLeibold, DCox, DAshby

Summary

Lost 20 minutes before opening fiddling with LDG Rotator.

Lost an hour because GCS wasn't configured to point to the machine (107) where the AGW4 version of oacserver was running. This is issue #2465.

Lost 2.25 hours to high winds. Spent alot of that time trying to debug a message that says: rejected change of guidestar list as new list is empty. This problem seems to be intermittent, and only happens to Doug.

Lost 1.5 hours trying to find the first star, and then discovering that the tertiary mirror was still deployed.

Lost 2 hours (still working on sky) convincing ourselves that the AGW4 guide CCD is broken. The saturation level is 3x lower than before, and we can't detect star images although it sees light. This is issue #2466. This caused us to switch back to AGW3 work at LFBG for the rest of the night.

The one bit of productive news is that the LDG collimation is still good as it was on Nov 25-27.

We also demonstrated that we can efficiently switch from LDG to LFBG in 20 minutes during the night.

Lost 1.4 hours because a DSP configuration file wouldn't let us switch back to the LFBG rotator. This is issue #2467.

We also had a problem that GCS 2.2.8B4 wouldn't display ACQUIRE images coming from PEPSI2 in AGW3, so we had to rollback to GCS 2.2.8B3. This is issue #2468.

Things settled down toward the end of the night and we got some productive work done. We've adjusted out the LFBG LUCIFER collimation and pointing problems.

Doug took some field aberration data on BS9143 in mediocre seeing. Then we uncovered another mystery about why Doug's IDL code isn't converging (sometimes) at LFBG.

Preparations

Jesper Storm wrote:
I have left a revised configuration file for AGW4 in

/lbt/oacontrol/etc/oacontrol.conf.20091223

which is based on the dataset from Oct.27. The fitting RMS is about 45
microns in x and y, so it seems very clean.
You will have to replace the oacontrol.conf file in that directory with the
new file and restart the AGW4 for the changes to take effect.

(We installed this configuration file later in the night, but we never got to test it because of CCD problems.)

Details

All Times in UT

00:20 Dome was parked at Azimuth 60 deg before sunset. Sky has scattered cirrus, with heavier cirrus moving in from the west. These clouds were enough to make a good sundog as the sun went down. Vent doors were open to cool things off.

00:40 David has succeeded in coercing the LDG rotator into functioning. It seems to have a similar problem as the LFBG rotator. Lost some pre-opening preparation time to this.

Open

00:40 T=-6.2 degC D=-31 degC Wind is slowly increasing over the last hour toward peak gusts of 20 m/sec.

The polycom unit on the cart has only blue screen video despite several attempts at power cycling.

01:03 Doug sends first preset. Fails because David had AGW3 turned on and AGW4 was off. But then problems with starting AGW4. (See Issue #2465, and continuing notes below.)

01:18 Messed around for a while with takepic until I discovered that I had IRTC.sendcam turned off.

01:24 We've discovered that the startup script for check4nports is not starting the daemon for agw4 (consistent with the error message we get about General Error). Calling Torsten.

Aaaargh - John accidentally deleted all the IRAF scripts from /home/LBTO/iraf/. (mostly a problem of oxygen deprivation) The following files were lost. Some important, some not. At least, all but one or two of them were John's own scripts.

-rw-r--r--  1 LBTO observers      656 Apr 15  2008 inex.cl
-rw-r--r--  1 LBTO observers     2133 Apr 17  2008 mkpupil.cl
-rw-r--r--  1 LBTO observers     1145 Apr 17  2008 spiral.cl
-rw-r--r--  1 LBTO observers     1145 Apr 18  2008 spiral25.cl
-rw-r--r--  1 LBTO observers      913 Apr 19  2008 inex_seq.old.cl
-rw-r--r--  1 LBTO observers     3218 Jun 21  2008 login.cl.orig
-rw-r--r--  1 LBTO observers     1102 Nov 15  2008 inex_seq.cl
-rw-r--r--  1 LBTO observers      360 Dec 22  2008 ccc.cl
-rw-r--r--  1 LBTO observers     4101 May 12  2009 fieldstar.cl
-rw-r--r--  1 LBTO observers     1806 May 12  2009 field_StoneM.cl
-rw-r--r--  1 LBTO observers      279 May 13  2009 tran_BS9152_90.cl
-rw-r--r--  1 LBTO observers      231 May 13  2009 tran_BS9152_00.cl
-rw-r--r--  1 LBTO observers      250 May 13  2009 tran_BS9152_180.cl
-rw-r--r--  1 LBTO observers      347 May 13  2009 tran_StoneL_00.cl
-rw-r--r--  1 LBTO observers      378 May 13  2009 tran_StoneM_00.cl
-rw-r--r--  1 LBTO observers      478 May 13  2009 tran_StoneM_90.cl
-rw-r--r--  1 LBTO observers      551 May 13  2009 tran_StoneM_180.cl
-rw-r--r--  1 LBTO observers      304 Jun 24  2009 tran_StoneO_00.cl
-rw-r--r--  1 LBTO observers      241 Sep 10 09:37 tran_StoneA_180.cl
-rw-r--r--  1 LBTO observers      283 Sep 10 09:38 tran_StoneA_90.cl
-rw-r--r--  1 LBTO observers      280 Sep 10 09:42 tran_BS9107_90.cl
-rw-r--r--  1 LBTO observers      351 Sep 10 09:47 tran_BS9107_180.cl
-rw-r--r--  1 LBTO observers      319 Sep 10 09:47 tran_BS9107_00.cl
-rw-r--r--  1 LBTO observers      221 Oct 10 06:01 tran_StoneA_270.cl
-rw-r--r--  1 LBTO observers      156 Oct 10 10:19 tran_StoneC_90.cl
-rw-r--r--  1 LBTO observers      185 Oct 10 10:26 tran_StoneC_180.cl
-rw-r--r--  1 LBTO observers     3355 Oct 11 04:32 transtar.20091011.cl
-rw-r--r--  1 LBTO observers    18543 Oct 12 03:50 mktpfile_new.cl
-rw-r--r--  1 LBTO observers      237 Oct 26 10:39 tran_StoneE_00.cl
-rw-r--r--  1 LBTO observers      282 Oct 26 10:46 tran_StoneE_180.cl
-rw-r--r--  1 LBTO observers      254 Oct 26 11:16 tran_StoneE_90.cl
-rw-r--r--  1 LBTO observers     5544 Oct 26 12:00 gopoint.20091026.cl
-rw-r--r--  1 LBTO observers      294 Oct 27 03:42 tran_StoneO_90.cl
-rw-r--r--  1 LBTO observers     3449 Oct 27 04:35 transtar.cl
-rw-r--r--  1 LBTO observers     5901 Oct 27 04:36 gopoint.cl
-rw-r--r--  1 LBTO observers      263 Oct 27 05:47 tran_StoneA_00.cl
-rw-r--r--  1 LBTO observers      293 Oct 27 06:59 tran_StoneO_270.cl
-rw-r--r--  1 LBTO observers      352 Oct 27 07:00 tran_StoneO_180.cl
-rw-r--r--  1 LBTO observers      252 Nov 25 08:01 tran_StoneC_190.cl
-rw-r--r--  1 LBTO observers      267 Nov 25 08:03 tran_StoneC_200.cl
-rw-r--r--  1 LBTO observers      293 Nov 25 08:04 tran_StoneC_225.cl
-rw-r--r--  1 LBTO observers      303 Nov 25 08:07 tran_StoneC_270.cl
-rw-r--r--  1 LBTO observers      251 Nov 25 08:09 tran_StoneC_315.cl
-rw-r--r--  1 LBTO observers      236 Nov 25 08:10 tran_StoneC_00.cl
-rw-r--r--  1 LBTO observers      279 Nov 25 11:06 tran_StoneE_270.cl
-rw-r--r--  1 LBTO observers     1931 Nov 27 04:19 field_BS9126.cl
-rw-r--r--  1 LBTO observers     1785 Nov 27 04:27 field_BS9152.cl
-rw-r--r--  1 LBTO observers     3648 Nov 27 05:05 field_BS9107.cl
-rw-r--r--  1 LBTO observers     1318 Dec  5 03:39 field_BS9101.cl
-rw-r--r--  1 LBTO observers      756 Dec  5 05:28 field_BS9101b.cl
-rw-r--r--  1 LBTO observers     3440 Dec 13 00:07 login.cl

The relatively important scripts gopoint.cl and transtar.cl I had saved downtown. And there was a backup copy of login.cl although it won't have the last changes made on Dec 13. The content of the tran_*.cl files can be recovered from the wiki logs with alot of typing. See Issue #2464 for my data recovery request. Stephen restored all of these files on 29-Dec.

Changing GCS configuration files for this

02:05 Torsten called to report that lbtmu105 (and 104) were not at all configured correctly to use AGW4. Torsten was all bent out of shape by this. But Torsten hadn't read the Thomas Hahn email of Dec 4. So we certainly have people communication problems that are contributing to this technical issue.

Close

02:12 Close for wind gusts greater than 20 m/sec. The sky is improved to only thin cirrus.

02:12 Changed oaccontrol_IP from agw-control to lbtmu107 on /lbt/tcs/current/etc/GCS/AIP_IRTC_DG_L.cfg

02:16 Stop and restart GCS

02:20 startAGW -u 4 (on lbtmu107)

02:22 Test TRACK preset succeeds.

02:28 Doug having problems with the GUIDE preset (from IRAF)

Sun Dec 27 02:28:28.985 2009  Preset command started for left IRTC mode GUIDE from IIF
Sun Dec 27 02:28:30.537 2009  Preset alert for left IRTC mode GUIDE: OK
Sun Dec 27 02:28:30.587 2009  Preset alert for left IRTC mode GUIDE: OK
Sun Dec 27 02:28:30.637 2009  Preset alert for left IRTC mode GUIDE: rejected change of guidestar list as new list is empty
Sun Dec 27 02:28:30.686 2009  Preset command failed for left IRTC mode GUIDE from IIF

 Dec 27 02:30:50 tcs4 LBT_left_GCS: GCSL:setGuidestarList executed from IIF, handle 6
Dec 27 02:30:50 tcs4 LBT_left_GCS: rejected change of _guidestarList as new list is empty.
Dec 27 02:30:50 tcs4 LBT_left_GCS: gcs.command.setGuidestarList.parameter.bad: rejected change of guidestar list as new list is empty

02:40 Doug sending preset from IDL - This seems to give a legitimate failure.

Sun Dec 27 02:40:57.792 2009  Preset alert for left IRTC mode GUIDE: AGw command status: AGW4: Requested position is Out of Range
Sun Dec 27 02:40:57.843 2009  Preset alert for left IRTC mode GUIDE: error: setProbePosition: couldn''t move probe to position (SFP) (x,y,z): 0.0858499, 204.965, 12.9 mm.
Sun Dec 27 02:40:57.894 2009  Preset alert for left IRTC mode GUIDE: guidestar list successfully changed but probe couldn''t be moved to position.
Sun Dec 27 02:40:57.945 2009  Preset command failed for left IRTC mode GUIDE from IIF
Sun Dec 27 02:41:02.153 2009  setting none mode on left side
Sun Dec 27 02:41:44.990 2009  Preset command started for left IRTC mode GUIDE from IIF
Sun Dec 27 02:41:47.781 2009  Preset alert for left IRTC mode GUIDE: OK

02:54 IDL GUIDE preset seems to succeed on-axis (with dome closed) up to the point of finding a star.

02:58 Doug sending IRAF version of same preset. This also succeeds in moving the probe.

03:00 M5_0050 gs=1 GUIDE preset seems to send probe to the wrong on-axis position. ??? ???
Dec 27 03:01:08 tcs4 LBT_left_GCS: AGWUnitAIP::doSetProbePosition: calling oacontrol::setxy(lbtmu107, 4, x:0.653351 y:612.484137, 23.000000, 0) ...

03:03 pointto M5_0050 TelM="GUIDE" gs=2

03:06 pointto M5_0050 TelM="GUIDE" gs=1 - It seems that gs1 really is at too large a radius, and that the probe is correctly not wanting to go there. But then we can't explain the message from 03:01 where GCS seems to move the probe on-axis.

03:13 pointto M5_0050 TelM="GUIDE" gs=3 - Also seems to be out of range, but here it correctly didn't move probe.

03:14 pointto M5_0050 TelM="GUIDE" gs=6

03:17 Stop and restart GCSL to see if we can reproduce the error from 02:28.

03:18 Preset to M5_0050 gs=0 (on-axis)

Sun Dec 27 03:18:55.498 2009  Preset alert for left IRTC mode GUIDE: OK
Sun Dec 27 03:18:55.549 2009  Preset alert for left IRTC mode GUIDE: OK
Sun Dec 27 03:18:55.599 2009  Preset alert for left IRTC mode GUIDE: AGw command status: AGW4: Requested position is Out of Range
Sun Dec 27 03:18:55.649 2009  Preset alert for left IRTC mode GUIDE: error: setProbePosition: couldn''t move probe to position (SFP) (x,y,z): -31.053, -195.848, 14.2052 mm.
Sun Dec 27 03:18:55.699 2009  Preset alert for left IRTC mode GUIDE: guidestar list successfully changed but probe couldn''t be moved to position.
Sun Dec 27 03:18:55.748 2009  Preset command failed for left IRTC mode GUIDE from IIF

This same preset fails three times in a row on the third try by Doug with IRAF.

03:20 John sends the same preset: pointto M5_0050 TelM="GUIDE" gs=0 and it works.

03:23 Now it works for Doug too. Doug is on obs3 and John is on obs2.

03:24 Stop and restart GCSL

03:25 Doug sends preset to gs=0, and it succeeds on the first attempt. So the wierdness is not repeatable. ???

03:28 Stop and restart GCSL

03:29 Doug sends preset to gs=0, and it succeeds on the first attempt.

03:30 Authorize LFBG then back to direct.

03:31 Doug sends preset to gs=0, and it succeeds.

03:32 Stop IIF and restart, and reauthorize.

03:31 Doug sends preset to gs=0, and it fails.
Sun Dec 27 03:33:58.253 2009  Preset alert for left IRTC mode GUIDE: OK
Sun Dec 27 03:33:58.304 2009  Preset alert for left IRTC mode GUIDE: rejected change of guidestar list as new list is empty
Sun Dec 27 03:33:58.355 2009  Preset command failed for left IRTC mode GUIDE from IIF

He repeats 4 times with failures

03:36 John's sending of the same preset succeeds. pointto M5_0050 TelM="GUIDE" gs=0

Doug's IRAF pointto request

IRTC> pointto M5_0050 gs=0 TelMode=GUIDE

         LBTtools -->>  IRTC:pointto.cl  <<-- Sun 03:35:15 27-Dec-2009

Log file will be written to: LOGs/pointto_033515.log

Current working directory: /OldRepository/AGW_Data/091226_iraf

Using source catalog: IRTC$/Catalogs/LBTdefault.cat

Using telescope mode = GUIDE

Science target: M5_0050
   Years to correct for proper motions: 9.987899543379
   Coordinates: 1:31:29.0320 15:20:44.750 2000.0
   (in radians) 0.39917367 0.26783411 
   (in degrees) 22.87096657 15.34576391 
   
Guide star: GS00
   Using science target as guide star!
   Coordinates: 1:31:29.0320 15:20:44.750 2000.0  R2mag=5.
   Offset: M5_0050->GS00 = 0.00 arcsec (0.00 W 0.00 S)
   (in radians) 0.00000000 0.00000000 

AGw Information:
   Set rotator to PA = 0. for AGw theta=0
   For PA = 0.00 -----------------------------
      PCS_XY coords:    0.000    0.000
      AGw_XY coords:    0.000  612.500

Writing the target.dat file
   # Target coordinates for source M5_0050
   1:31:29.0320 15:20:44.750 RADEC 2000.0 J2000 0.0 0.0 5.0 R 0.0 H_K 1.4
   # Guide Star coordinates for GS00
   1:31:29.0320 15:20:44.750 RADEC 2000.0 J2000 0.0 0.0 5. R 0. B_V 0.63

   #############################################################
   ###  Sending the following command sequence to the IRTC:  ###
   #############################################################
   irc ClearStars
   irc SetHotspot 0.0 0.0
   irc SetOffset 0.0 0.0 RADEC
   irc SetStars target.dat
   irc PresetTelescope 0. POSITION GUIDE left
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = OK
Resmsg[2]  = OK
Resmsg[3]  = rejected change of guidestar list as new list is empty

   to examine the log: page LOGs/pointto_033515.log

03:43 Doug's preset fails after he's made a new proxy object for IIF with the register command.

03:45 John's preset still works after he makes a fresh proxy.

03:47 Doug's preset succeeds.

03:48 Doug registers as LUCIFER and the preset fails as it should

03:49 Doug registers as DGTEST (IRTC at LDG), and his preset works.

03:50 Slew around to East to check wind.

03:57 Copied Jesper's new oacontrol.conf.20091223 file into oacontrol.conf on /lbt/oacontrol/etc/
But we learn a few days later that this does not affect lbtmu107, but only lbtmu105/104.

03:58 Stop GCSL, stopAGW -u 4, stop oacserver, and restart them all.

04:07:41 Doug's preset to M5_0050 gs=0 works properly

04:29:09 David stops and restarts IIF, authorized IRTC at LDG

04:11:12 Doug's preset to M5_0050 gs=0 works properly. There does not seem to be pattern to the "empty list" nor to the incorrect probe position (should have been on-axis) problems we have seen from my Doug's IRAF

Open

04:27 Open dome as the wind has decreased to 12-15 m/sec.

04:32 Preset to M5_0256 gs=0 ACQUIRE - ptspiral, but no star found. Notice that there was a 1.2 mm secondary offload. Where did it come from? (Changing authorizations?) We cleared it. But we still can't find a star.

04:58 We point in the vicinity of the moon, and we verify that the AGW is seeing photons.

05:00 Go back to 120 deg Azimuth to verify the encoder reset position. David reset it again.

05:05 pointto M5_0258 ACQUIRE GS=0

05:12 Check az encoder again - it was within 1 arcsec at 120 deg.

05:15 pointto Pollux gs=0 ACQUIRE - and we aren't close.

Verified that we are talking to the correct AZCAM by stopping AGW3 (had to kill oacserver before this could be done).

05:25 stop GCSL, stopAGW -u 4 (to reset everything) and home probe again.

05:35 Found that Tertiary Mirror is still in the beam !!!!

05:38 Go to zenith to remove tertiary mirror and adjust dynamic balance.

06:00 pointo M5_0258 ACQUIRE gs=0 - still don't see star. So we ptspiral again.

06:06 found star at *IE=0 CA=+122 * but it is very strange looking. Seems to have some god-awful amount of coma. Or else the CCD is broken...........

06:14 Send TRACK mode preset so we can look at the wierd image on IRTC. The wierd image looks OK on IRTC, no coma and only a little bit out of focus. Refined *IE=7 CA=115 * .

06:20 We can't see off-axis guide star 7

06:35 Move to gs=5, can't see it on the acquire movie either. There is a lovely 1 arcsec image on IRTC. This confirms that LDG collimation model is still quite good.

Can something have fallen onto the AGW probe to block the image?

06:42 Doug presets us to the moon again. Guider saturates on the moon in a millisecond, but still the image looks very strange on the CCD. see /tmp/moon1.fits, moon2.fits, moon3.fits as the limb drifts out of the frame. These images got moved to /OldRepository/AGW_Data/091227/.

Close

07:02 Close dome and go to zenith to inspect AGW probe. Inspection from what you can see leaning over the edge of the AGW doesn't show any debris. (And its cleaner up there than I was expecting.) So the soft conclusion is that the guide camera has a bad CCD or a bad cable or a bad power supply.

Switch back to LFBG (seems like we were just there)

07:30 Change configuration back to LFBG. (deploy M3 and rebalance)

Taking some test data with AGW4 guide CCD:

in /OldRepository/AGW_Data/091227/

guider0001 1 sec "dome" flat (back of M3)
guider0002-4 1 sec "dome" flat with telescope at horizon
guider0005-7 10 sec
guider0008-9 20 sec
guider0010-11 5 sec
guider0012-13 2 sec
lights off
guider0014-15 2 sec with lights off
guider0016 20 sec with lights off

Open

07:54 Re-open shutter doors and authorize as LUCIFER.

LFBG rotator problem (feature)

08:00 Problem with LFBG rotator. Restart MCSPU.
Dec 27 07:45:30.300: MCSPU: rhold called with arg = 'LFBG'.
Dec 27 07:45:30.300: MCSPU: rhold result = BAD rotator. This rotator doesn't exist.

08:20 David's operation of the MCR reset button gets the LDG rotator working fine, but the same cure does not work for the LFBG rotator.

08:29 reboot jet - didn't help

08:41 call Dan Cox, but he doesn't answer.

[telescope@jet mcspuLogs]$ grep LFBG 091* |grep exist
091216.2005:Dec 16 20:05:37.434: LFBG Rotator: Rotator does not exist!
091226.1612:Dec 26 16:12:44.848: LFBG Rotator: Rotator does not exist!
091226.2246:Dec 26 22:46:51.874: LFBG Rotator: Rotator does not exist!
091226.2349:Dec 26 23:49:14.840: LFBG Rotator: Rotator does not exist!
091227.0019:Dec 27 00:19:17.202: LFBG Rotator: Rotator does not exist!
091227.0026:Dec 27 00:26:27.195: LFBG Rotator: Rotator does not exist!
091227.0803:Dec 27 08:03:26.471: LFBG Rotator: Rotator does not exist!
091227.0836:Dec 27 08:36:43.270: LFBG Rotator: Rotator does not exist!

08:55 call Dave Ashby. He sorts out that the LFBG rotator DSPs are left out of the configuration file.

09:21 Dave sorts out a configuration file that will run LFBG but not LDG (what we need for the rest of tonight).

09:30 pointto M5_0314 gs=0 ACQUIRE - GCSGUI won't display a valid image even though AZCAM seems to be taking one every 5 seconds.

Restart GCSGUI - didn't help

09:39 Restart GCSL. resend preset. No help.

09:41 Stop GCSL, stopAGW -u 3, restart, home, etc. in order to reboot AzCam.

The wierd thing I seem in the image header of movie.fits is that DARKTIME is -842150.451.

09:46 pointto M5_0314 gs=0 ACQUIRE - GCSGUI still won't display image.

09:49 pointto M5_0314 gs=0 GUIDE - This displayed the acquisition image.

09:58 Stop and restart GCSL to roll back to V2.2.8 B3.

10:00 pointto M5_0314 gs=0 ACQUIRE -- finally we see a star image of sorts (modulo focus and coma) on the guide camera. This means that GCS V2.2.8 B4 has a bug that doesn't allow it to display the acquisition movie images from PEPSI2=AGW3

10:02 resent preset in parallactic mode

Fixing Up the LFBG pointing and collimation

10:06 Moved M2 to the Direct Position this moved it from Y=+2 to Y=-1.3.

10:10 Dan C called to tell us that the configuration file change was deliberate because of a problem in telemetry when both rotators are sending telemetry. Life is never simple.

10:16 Put M2 back to the LFBG default (but M2 didn't move when I did that - did I need to hit collimate after initialize?). And move M1 to the LDG angle. This gets rid of the horrible coma.

M1 Global is X=-0.583 Y=-2.633 Z=0.181 RX=-39.2 RY=+4.87
M2 Global is X=-7.010 Y=+2.121 Z=0.000 RX=+148.65 RY=+181.00

10:20 When we really put M2 at the LFBG position......then the coma is back with IE=-78 CA=13.

10:30 M2 Y=-0.3 wasn't enough, then -1.3 reversed the coma, so we adopt the compromise:

10:32 M2 Y=-0.6 mm

M1 Global is X=-0.583 Y=-2.633 Z=0.181 RX=-39.2 RY=+4.87
M2 Global is X=-7.010 Y=-0.600 Z=0.000 RX=+148.65 RY=+181.00

Seeing is about 1 arcsec on the guider after a bit of Active Optics. IE=-19 CA=+13 at EL=70

10:40 focus 0 to force offload

Accounting for the offload plus global combined, the M1 global should be:
M1 Global is X=-0.980 Y=-2.216 Z=0.403 RX=-34.78 RY=+9.0
at EL=70.3

10:47 pointto M5_0354 gs=0 TelM="ACTIVE" (That star didn't seem to have any off-axis guide stars.)

10:51 pointto M5_0354 gs=0 TelM="ACQUIRE" PAM- IE=-15 CA=+25 at EL=48 AZ=191.

10:55 pointto M5_0419 gs=0 TelM="ACQUIRE" PAM- *IE=-13 CA=31 * at EL=28 AZ=163 with IA=-290

11:03 Make IA=-260 CA=+3 IE=-13. As we learned on Nov 26, IA and CA sum together so to make CA more negative we need to make IA more positive.

11:04 Slew to 120 deg to check the AZ zeropoint. This made a shift of 1.5 arcsec.

11:09 pointto BS9143 gs=0 TelM="ACQUIRE" PAM- Adjust as IE=-14 CA=5.5 IA=-260 at EL=61.7 AZ=170.

11:13 pointto BS9143 gs=0 TelM="ACTIVE" - Seeing is running 1.1 to 1.4 arcsec, so not very good for field data.

11:18 Adjusted LUCIFERSX.20091227.ptmod to have the new IA value of -260. Note that this change essentially puts us back to the 24-Sep-2009 version of the LUCIFER pointing model.

T   51   3.7508    0.000   0.0000
  IA       -260.0000     6.28333
  IE        -14.0000     0.93987
  NPAE       +6.0495     7.84478
  AW         -3.4341     0.60172
  AN        +22.5359     0.58791
  CA         +5.5000     9.77808
  TX        +13.1546     0.64348

11:20 Adjusted SXSMLUCIFERCollimation.dat to M2 Global is X=-7.010 Y=-0.600 Z=0.000 RX=+148.65 RY=+181.00

11:30 Adjusted SXPMLUCIFERCollimation.dat to M1 Global is X=-0.980 Y=-2.216 Z=0.403 RX=-34.78 RY=+9.00

LFBG Field Data

Data in /OldRepository/AGW_Data/091227_wfs

Seeing 1.0" to 1.5" (occasionally 2.0")

Note that this seeing, and these data are effected by the IDL non-convergence problem that we found below.

Persson's Star = BS9143
Start time = 11:38

Guide Star x y PA Doug's Angle wfscimage000*
0 0.22 612.62 0 352 2-3
18 -0.27 492.19 168.79 508 4-6
0 0.1 612.5 168. 508 7-8
26 3.7 435.9 80.13 426 9-11
0 -0.07 0.06 80.13 426 12-13
14 0.03 109.70 9.43 352 14-16
0     9.43 352 17-18
2 0.18 195.35 279.13 258 19-21
0 0.12 -0.03 279.13 256 22-23
8 -0.01 93.50 203.53 179 24-26
0 0.03 -0.17 203.53 537 27-29
22 -0.06 141.43 43.83 376 30-32
0 -0.01 0.17 43.83 376 33-34
11 -0.09 143.58 177.35 506. 35-37
0 -0.09 -0.16 177.35 506. 38-39
25 -0.11 168.62 58.53 385 40-42
0 -0.04 0.18 58.53 385 43-44

End time = 12:59

13:00 preset to BS9152 with
cl < home$newfield_BS9152.cl
to test a recovered field data script.

Working in /OldRepository/AGW_Data/091227/

But the agw_collimate isn't giving the right corrections, the image diverged. See wfsc0001-4.

13:27 agw_collimate, exp=30 by hand by John

13:31 agw_collimate, exp=30 by hand by Doug - both of these seem to be diverging the image. Yet Doug was able to converge a while ago.

Doug found and fixed this problem the following afternoon.

13:35 pointto BS9152 gs=0 ACTIVE - this converges OK to 1.0 arcsec image. Why does IIFGUI continuously show that WFS is paused?

Close

13:48 Close for impending dawn. Sky is clear. T=-6.9 degC D=-25.9 degC. Wind ~10 m/sec from West

-- JohnHill - 27 Dec 2009
Topic attachments
I Attachment Action Size Date Who Comment
daq.20091227.jpgjpg daq.20091227.jpg manage 789 K 28 Dec 2009 - 00:11 JohnHill Plot of Telescope and Mirror Temperatures
pmcc.dx.thermal.20091227.jpgjpg pmcc.dx.thermal.20091227.jpg manage 675 K 28 Dec 2009 - 00:14 JohnHill Plot of DX primary mirror tempertures
pmcc.sx.thermal.20091227.jpgjpg pmcc.sx.thermal.20091227.jpg manage 531 K 28 Dec 2009 - 00:13 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r11 - 30 Dec 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback